MONT 105Q – Mathematical Journeys First Writing Assignment due: February 26, 2015

The first of the larger paper assignments this semester will be devoted to our second CHQ common “text,” the film The Curious Case of Benjamin Button, directed by David Fincher and starring Brad Pitt and Cate Blanchett. Note: If you were unable to see the film at the CHQ showing on February 18, you may either watch the film on Netflix, or write your paper about the F. Scott Fitzgerald short story with the same title (see link on course homepage).

An interesting sidelight on the film and the short story is that F. Scott Fitzgerald says his story was inspired to some extent by a quotation attributed to Mark Twain:

Life would be infinitely happier if we could only be born at the age of eighty and gradually approach eighteen.

For this assignment, you will need to come up with your own topic connected with this general idea about how the direction of time impacts our lives. The following are some leading questions and suggestions to get you started. Don’t feel you need to address all (or any of) these points if you want to say something quite different, though.

• What is Twain trying to say in his statement? Note: it’s usually taken as humorous because of the way it subverts our expectations about the way time works. It was also apparently something he wrote in the final years of his life.

• To what extent does the film (or the short story) actually bear out Twain’s statement? Note that Benjamin’s life arc does not stop at eighteen! Is that life arc actually a happier one than the normal one in which humans age in the opposite direction? Or is the film (or the short story) perhaps saying that Twain had it all wrong?

• From your (necessarily limited) persective, is it possible to agree with Twain? Is where you are now essentially where you would want to be ending your life? Or is the rest (i.e. whatever happens between now and eighty) worth experiencing whatever pain and loss comes with it?

• Are there “life lessons” about aging that you can identify in the film (or the short story)? How does that relate to this year’s CHQ theme?

• In a different direction, if you want to do a “compare and contrast” paper, you might discuss the differences between the short story and the film and consider why the screenwriters chose to keep what they kept from the story and to change what they changed. Your writing should show that you have done some serious thinking about the movie

and/or the short story and the issues they raise. You may, of course, disagree with ideas presented in our class or in the discussion with the students from Professor Vineberg’s seminar if you want. The point as always is for you to develop your own thinking, and then clearly express it in writing.

Your essay should be 3-5 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Friday, February 26.

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. But of course your essay should be entirely your own work. I will also be happy to talk over things and/or read a preliminary draft to give some feedback if you like.

2. The first paragraph(s) should clearly lay out what topics and/or questions you will consider. This will form the thesis and it should be clearly stated. The reader (me) should not need to guess what your point is.

3. The paper as a whole should be well-organized and “flow” from one point to the next. Your sentences should be clear, readable, and not overly wordy.

4. This paper can be less formal. You should aim for a lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason.

5. If you feel you must, it will be OK to consult other sources for this assignment. If you do consult reviews of the film, for instance, please provide the author’s name, the title, and the name of the periodical or the full URL of the web page where the review appeared. For a web page, also provide the date you consulted it online. All sources you use should be included as a numbered list in a section called References at the end of your paper. These can be ordered according to where they are cited in the text if you prefer.

5. If you use direct quotations, include footnotes to indicate the source, which then must appear in your References section.

6. Please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

7. Errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.